A store has wrapping paper on sale. One package is 3 rolls of 2 1/2 ft x 8 ft for $6.00. Another package is 5 rolls of 2 1/2 ft x 6 ft for $8.00. Which is the better buy per square foot?
Answer by stanbon(75887)

A store has wrapping paper on sale.
One package is 3 rolls of 2 1/2 ft x 8 ft for $6.00.
A: Price per square foot: 600/[3*(5/2)*8] = 600/60 = 10 cent per sq. ft.
---------
Another package is 5 rolls of 2 1/2 ft x 6 ft for $8.00.
B: Price per sq. ft: 800/[5*(5/2)*6] = 10.67 cents per sq ft.
------------------------------------------------
Which is the better buy per square foot?
A is the better buy
==========================================
Cheers,
Stan H.
===================